(a) Two parallel, very long wires 1 and 2 are separated by a distan
e of 0.065 m and
arry
urrents
of I1 = 15 A and I2 = 7 A. Find the magnitude of the for
e that the magneti
eld that wire 1
applies to a 1.5 m of wire 2 when the
urrents are in opposite dire
tion.
(b) A 150-turn solenoid
arries a
urrent of 12 A. The radius of the solenoid is 0.050 m; and its length
is 0.18 m. Determine the magneti
ux through the
ir
ular
ross-se
tional area at the
enter of
the solenoid.
1
Expert's answer
2020-05-04T12:40:42-0400

a) As per the given question,

There is two long parallel wire,

Current flowing in the two wire in the opposite direction, that is "(i_1)=15A" and "(i_2)=7A"

distance between the wire is "(d)=0.065m"

magnetic field from the wire 2, at a distance 1.5

"B_{net}=B_1-B_2=\dfrac{\mu_o}{4\pi}(\dfrac{i_2}{R_2}-\dfrac{i_1}{d+R_1})"


"\Rightarrow B=\dfrac{\mu_o}{4\pi}(\dfrac{7}{1.5}-\dfrac{15}{1.565})=10^{-7}\times(\dfrac{-11.545}{2.3475})"


"\Rightarrow B=4.917 \times 10^{-7}T\simeq 5.0\times 10^{-7}T"